# sixth generation AE autoencoder

### v1

@@ V-B3 Low-complexity LLR Calculation
We observe that the LLRs of the first and second bit depend mainly on the amplitude (and not the phase) of the received symbol. Therefore, we transform these LLRs $l_{1}(\operatorname{Re}\{y_{\text{c}}\},\operatorname{Im}\{y_{\text{c}}\})$ and $l_{2}(\operatorname{Re}\{y_{\text{c}}\},\operatorname{Im}\{y_{\text{c}}\})$ from Cartesian into polar coordinates, where $R=|y_{\text{c}}|$ and $\theta=\sphericalangle\{y_{\text{c}}\}$. Then, we average over the angular dimension $\theta$
@@
Fig. 7 shows the marginalized LLR functions for all six bits, which can now be approximated using one-dimensional LUTs, enabling low-complexity implementation. Interestingly, the radial components of bits 5 and 6 are identical, such that a single LUT suffices for both. Therefore, only seven 1D look-up tables are needed in total: one each for bits 1-4, one for the radial component shared by bits 5 and 6, and one for their respective phase components. We note that the one-dimensional LLRs can be approximated using piecewise linear functions, further reducing complexity. However, the specific implementation details are beyond the scope of this paper. Nevertheless, assuming $V$ entries per dimension per LUT, this reduces the total storage requirement from $\mathcal{O}(6V^{2})$ to $\mathcal{O}(7V)$, offering a significant memory saving. We remark that the proposed approximation generalizes to larger constellations: Amplitude bits remain largely insensitive to phase, the first two phase bits correspond to real and imaginary half-planes, and higher-order phase bits exhibit oscillatory angular patterns.
@@ VI-B Sensing Performance
In Fig. 9, we consider a scenario with two targets: a distant interfering target and a nearby TOI, which should be detected and whose distance is varied. Multi-target scenarios are of particular interest because the average SINR (14) depends on the kurtosis and the power reflected by all targets. The simulation setup follows [6], with the simulation parameters being the FR2 case from [29] with only one OFDM symbol. For CA-CFAR, we assume a false alarm rate of $P_{\text{FA}}=10^{-3}$ and a sliding window length of $N_{\text{win}}=$100$\$.
We found that the simulated detection rates (markers) align well with the analytical (10) constellation-dependent detection probabilities (curves), verifying our derivation in Sec. II. The minor discrepancies stem from the fact that the kurtosis $\kappa$ of the optimized constellations slightly differs from the kurtosis constraint $\tilde{\kappa}$. As expected from (10) and (14), the detection probability decreases with increasing kurtosis $\kappa$ and depends only on the kurtosis of the constellation, irrespective of the specific shaping method. Moreover, our simulations demonstrate that constellation shaping enables a dynamic adjustment of the detection range. For example, for a required target detection probability of $P_{\mathrm{D}}=$0.9$\$, the detection range can be varied from $90\text{\,}\mathrm{m}$ to beyond $160\text{\,}\mathrm{m}$ by modifying the kurtosis $\kappa$ of the constellation.
@@ VI-C Communications Performance
Fig. 10 shows the GMI [23, Eq. 14] as a function of the communications SNR for various sensing constraints $\tilde{\kappa}$. For $\tilde{\kappa}=2$, all shaping methods reduce the gap to capacity and outperform conventional $64$-QAM across an SNR range of $10\text{\,}\mathrm{d}\mathrm{B}$. We note that probabilistic and joint shaping reduce the entropy of the transmit symbols at the design SNR of $\mathrm{SNR}_{\text{c}}=$10\text{\,}\mathrm{d}\mathrm{B}$\$, resulting in a saturation of the GMI below $6\text{\,}\mathrm{b}\mathrm{i}\mathrm{t}\mathrm{/}\mathrm{s}\mathrm{y}\mathrm{m}\mathrm{b}\mathrm{o}\mathrm{l}$ for large SNR values. For $\tilde{\kappa}=1$, the GMI of the shaped constellations is similar to that of the PSK, although probabilistic constellation shaping performs slightly worse due to unequal distances between the constellation points and the absence of Gray coding. Both geometric and joint constellations slightly outperform the $64$-PSK because their resulting constellations resemble a $16$-PSK which achieves a higher GMI at low SNR values.
@@ VI-D Sensing & Communications Trade-off
Next, we observe that the constellations optimized for SMD yield a MI that closely approaches the upper bound (18). In Fig. 13, we compare the empirical PDF of the received signal resulting from the optimized constellations to the maximum-entropy PDF used to derive the upper bound (20). For $\kappa=2$, both PDFs converge to the circularly symmetric Gaussian PDF, which achieves the capacity for an AWGN channel. At lower kurtosis values, the two PDFs diverge slightly, as the upper bound does not constrain the output to be realizable under the system model, resulting in a gap between the achieved MI and the derived bound (18). Nevertheless, the gap in Fig. 12 is less than $0.04\text{\,}\mathrm{b}\mathrm{i}\mathrm{t}\mathrm{/}\mathrm{s}\mathrm{y}\mathrm{m}\mathrm{b}\mathrm{o}\mathrm{l}$ at $\kappa=1$ and below $0.01\text{\,}\mathrm{b}\mathrm{i}\mathrm{t}\mathrm{/}\mathrm{s}\mathrm{y}\mathrm{m}\mathrm{b}\mathrm{o}\mathrm{l}$ for $\kappa\geqslant 1.05$, demonstrating both the tightness of the bound and the effectiveness of the AE-based constellation optimization.
@@
Figure 13: Comparison of the empirical received signal PDF resulting from the optimized constellations and the theoretical maximum-entropy PDF used to derive the upper bound for various kurtosis constraints, along with a Gaussian as reference. Note that the original PDFs are two-dimensional. The displayed PDFs are obtained by averaging one-dimensional marginals computed along 6464 uniformly spaced radial directions in the range $0,\pi]$.
@@
Remark: This makes geometric and probabilistic constellation shaping well-suited for applications where sensing or communications performance is prioritized, respectively. Joint shaping consistently exhibits strong performance across all regimes, effectively leveraging the advantages of both geometric and probabilistic shaping. It performs similarly to geometric and probabilistic constellation shaping at low and high kurtosis values, respectively. Compared to a conventional $64$-QAM, joint constellation shaping increases the GMI by 0.16 bit/symbol and reduces the kurtosis $\kappa$ by $0.19$ while maintaining an equivalent S&C performance. Additionally, while $64$-PSK and $64$-QAM offer only two discrete operating points, joint constellation shaping allows for continuous adjustment between S&C performance, providing greater flexibility.
@@ VII Conclusion
$=\pi\mathrm{e}^{\gamma_{0}}\mathrm{e}^{\frac{\gamma^{2}_{2}}{4\gamma_{4}}}\!\!\!\!\!\!\!\!\!\!\int\limits_{-\gamma_{2}/(2|\gamma_{4}|)}^{\infty}\!\!\!\!\!\!\left(v+\frac{\gamma_{2}}{2|\gamma_{4}|}\right)^{q}\mathrm{e}^{-|\gamma_{4}|v^{2}}\,\mathrm{d}v$ (43)
@@
Using the integrals [30, entries: 3.321.2, 3.321.4, 3.321.5], we obtain closed-form expressions
@@
$\begin{split}C_{2}&=\pi\mathrm{e}^{\gamma_{0}}\left[\frac{\gamma_{2}}{4|\gamma_{4}|^{2}}+\left(\frac{1}{4|\gamma_{4}|}+\frac{\gamma_{2}^{2}}{8|\gamma_{4}|^{2}}\right)\right.\\ &\quad\qquad\quad\cdot\left.\sqrt{\frac{\pi}{|\gamma_{4}|}}\mathrm{erfc}\!\left(\frac{-\gamma_{2}}{2\sqrt{|\gamma_{4}|}}\right)\mathrm{e}^{\frac{\gamma_{2}^{2}}{4|\gamma_{4}|}}\right].\end{split}\vskip-1.42271pt$ (46)
Using $C_{0}=1$ (Tab. I), we substitute the common term in (44) into (45) and (46) yielding
@@
$C_{2}$ $=\frac{\gamma_{2}}{4|\gamma_{4}|^{2}}\left(\pi\mathrm{e}^{\gamma_{0}}+\gamma_{2}\right)+\frac{1}{2|\gamma_{4}|}.\vskip-1.42271pt$ (48)

### v2

@@ V-B3 Low-complexity LLR Calculation
For convenience, we recall the LLR of bit position $m$, see 2
$l_{n,m}(y_{\text{c},n})\!=\!\log\frac{\sum_{x_{\!n}\in\mathcal{X}^{(0)}_{m}}{\!f_{\mathsf{y}_{\text{c}}|\mathsf{x}}(y_{\text{c},n}|x_{n})}P_{\mathsf{x}}(x_{n})}{\sum_{x_{n}\in\mathcal{X}^{(1)}_{m}}{\!f_{\mathsf{y}_{\text{c}}|\mathsf{x}}(y_{\text{c},n}|x_{n})}P_{\mathsf{x}}(x_{n})},$ (40)
which indicates whether bit $m$ is more likely to be 0 or $1$ and quantifies the reliability of this decision. We observe that the LLRs of the first and second bit depend mainly on the amplitude (and not the phase) of the received symbol. Therefore, we transform these LLRs $l_{1}(\operatorname{Re}\{y_{\text{c}}\},\operatorname{Im}\{y_{\text{c}}\})$ and $l_{2}(\operatorname{Re}\{y_{\text{c}}\},\operatorname{Im}\{y_{\text{c}}\})$ from Cartesian into polar coordinates, where $R=|y_{\text{c}}|$ and $\theta=\sphericalangle\{y_{\text{c}}\}$. Then, we average over the angular dimension $\theta$
@@
Fig. 8 shows the marginalized LLR functions for all six bits, which can now be approximated using one-dimensional LUTs, enabling low-complexity implementation. In the top-right plot, we observe that bits 3 and 4 depend on their respective Cartesian component in an equivalent manner. This follows from the rotational symmetry of the shaped constellation (see Fig. 7) and the fact that they represent coarse phase information, corresponding to the sign of the real and imaginary parts, respectively. Consequently, a single LUT suffices for both bits. Furthermore, the radial components of bits 5 and 6 are similar (see left plot) because both bits encode further phase information. Consequently, a single LUT suffices for the radial part of both bits. Therefore, only six one-dimensional look-up tables are needed in total: one each for bits 1 and 2, one for bits 3 and 4, one for the radial component shared by bits 5 and 6, and one for their respective phase components. We note that the one-dimensional LLRs can be approximated using piecewise linear functions, further reducing complexity. However, the specific implementation details are beyond the scope of this paper. Nevertheless, assuming $V$ entries per dimension per LUT, this reduces the total storage requirement from $\mathcal{O}(6V^{2})$ to $\mathcal{O}(7V)$, offering a significant memory saving. We remark that the proposed approximation generalizes to larger constellations: Amplitude bits remain largely insensitive to phase, the first two phase bits correspond to real and imaginary half-planes, and higher-order phase bits exhibit oscillatory angular patterns.
@@ VI-B Sensing Performance
In Fig. 10, we consider a scenario with two targets: a distant interfering target and a nearby TOI, which should be detected and whose distance is varied. The RCS of the TOI follows a Swerling-1 model, whereas the interfering target follows a non-fluctuating Swerling-0 model. Multi-target scenarios are of particular interest because the average SINR (21) depends on the kurtosis and the power reflected by all targets. To verify the impact of the constellation on the detection performance, we evaluate the resulting detection rate for each of the optimized transmit constellations. The simulation setup follows [6], with the simulation parameters being the FR2 case from [32] with N=12 672N=\$12\,672\$ sub-carriers. For simplicity, we use an FFT size of 12 67212\,672, and only one OFDM symbol. For CA-CFAR, we assume a false alarm rate of $P_{\text{FA}}=10^{-3}$ and a sliding window length of $N_{\text{win}}=$100$\$. To obtain the detection probability, we perform a Monte-Carlo simulation with $100$ independent Swerling-1 channel realizations. This inherently averages over the RCS fluctuations of the TOI, such that the mean target power converges to that determined by the mean RCS, see (21). Hence, the impact of the Swerling-1 fluctuations vanishes, and the detection probability is governed by the noise-and-interference power, which depends on the kurtosis $\kappa$ of the constellation, as described in Sec. II-E.
@@
We found that the simulated detection rates (markers) align well with the analytical (12) constellation-dependent detection probabilities (curves). This verifies our derivation in Sec. II and shows the impact of the kurtosis $\kappa$ of the constellation on the detection probability. The minor discrepancies stem from the fact that the kurtosis $\kappa$ of the optimized constellations slightly differs from the kurtosis constraint $\tilde{\kappa}$. As expected from (12) and (21), the detection probability decreases with increasing kurtosis $\kappa$ and depends only on the kurtosis of the constellation, irrespective of the specific shaping method. Moreover, our simulations demonstrate that constellation shaping enables a dynamic adjustment of the detection range. For example, for a required target detection probability of $P_{\mathrm{D}}=$0.9$\$, the detection range can be varied from $90\text{\,}\mathrm{m}$ to beyond $160\text{\,}\mathrm{m}$ by modifying the kurtosis $\kappa$ of the constellation.
@@ VI-C Communications Performance
[…]
@@ VI-D Sensing & Communications Trade-off
Next, we observe that the constellations optimized for SMD yield a MI that closely approaches the upper bound (25). In Fig. 14, we compare the empirical PDF of the received signal resulting from the optimized constellations to the maximum-entropy PDF used to derive the upper bound (27). For $\kappa=2$, both PDFs converge to the circularly symmetric Gaussian PDF, which achieves the capacity for an AWGN channel. At lower kurtosis values, the two PDFs diverge slightly, as the upper bound does not constrain the output to be realizable under the system model, resulting in a gap between the achieved MI and the derived bound (25). Nevertheless, the gap in Fig. 13 is less than $0.04\text{\,}\mathrm{b}\mathrm{i}\mathrm{t}\mathrm{/}\mathrm{s}\mathrm{y}\mathrm{m}\mathrm{b}\mathrm{o}\mathrm{l}$ at $\kappa=1$ and below $0.01\text{\,}\mathrm{b}\mathrm{i}\mathrm{t}\mathrm{/}\mathrm{s}\mathrm{y}\mathrm{m}\mathrm{b}\mathrm{o}\mathrm{l}$ for $\kappa\geqslant 1.05$. This demonstrates the effectiveness of the AE-based constellation optimization and shows that the upper bound can be closely approached by constellation shaping.
Figure 14: Comparison of the empirical received signal PDF resulting from the optimized constellations and the theoretical maximum-entropy PDF used to derive the upper bound for various kurtosis constraints, along with a Gaussian as reference. Note that the original PDFs are two-dimensional. The displayed PDFs are obtained by averaging one-dimensional marginals computed along $64$ uniformly spaced radial directions in the range $0,\pi]$.
@@
Remark: This makes geometric and probabilistic constellation shaping well-suited for applications where sensing or communications performance is prioritized, respectively. Joint shaping consistently exhibits strong performance across all regimes, effectively leveraging the advantages of both geometric and probabilistic shaping. It performs similarly to geometric and probabilistic constellation shaping at low and high kurtosis values, respectively. Compared to conventional $64$-QAM, joint constellation shaping achieves a $0.16$ bit/symbol higher GMI at the same kurtosis $\kappa$. This means that joint shaping increases the communication throughput while maintaining the same detection probability for sensing. Conversely, joint shaping can reduce the kurtosis $\kappa$ by $0.19$ while achieving the same GMI, which directly translates into a higher detection probability of the TOI at equal communication performance. In addition, while $64$-PSK and $64$-QAM offer only two discrete operating points, joint shaping enables a continuous adjustment of the AIR-kurtosis operating point. This allows for a flexible and fine-grained trade-off between communication throughput and sensing detection probability. Furthermore, remember that the self-interference from the limited isolation between the transmitter and receiver paths manifests as a target with a large virtual RCS, and that the noise-and-interference (16) floor scales with the product of the kurtosis and the summed power of all targets. Consequently, if the self-interference is stronger than the received reflections of the targets, it dominates the noise-and-interference floor, which scales with the kurtosis of the employed constellations. Hence, reducing the kurtosis through constellation shaping can reduce the noise-and-interference floor caused by the self-interference and, consequently, increase the detection probability of a TOI with small RCS.
@@ VII Conclusion
Future research includes extending the proposed framework to frequency-selective communication channels, where the communications SNR varies across subcarriers and power allocation becomes relevant, as discussed in Sec. II. Additional directions include incorporating the Doppler effect, i.e., velocity estimation, for sensing and studying constellation shaping under practical hardware impairments.
@@
To show that the kurtosis $\kappa$ of a conventional $\widetilde{M}-$QAM approaches $\kappa\approx 1.4$ as $\widetilde{M}\to\infty$ (see Sec. II-E), we use that, in this limit, the constellation converges to an independently and uniformly distributed random variable on $\mathcal{U}[-\sqrt{3/2},\sqrt{3/2}]$ in both dimensions, i.e.,
$\mathsf{x}=\mathsf{x}_{\text{Re}}\!+\!\mathrm{j}\mathsf{x}_{\text{Im}},\!\!\!\quad\mathsf{x}_{\text{Re}},\mathsf{x}_{\text{Im}}\!\sim\!\mathcal{U}\big[\!-\!\sqrt{3/2},\sqrt{3/2}\big],\!\!\!\quad\mathsf{x}_{\text{Re}}\!\perp\!\mathsf{x}_{\text{Im}}.\vskip-5.69046pt$ (47)
The kurtosis then directly follows as
$\mathbb{E}\{|\mathsf{x}|^{4}\}$ $=\!\mathbb{E}\!\left\{\!\left(\mathsf{x}^{2}_{\text{Re}}\!+\!\mathsf{x}^{2}_{\text{Im}}\right)^{\!2}\!\right\}\!=\!2\!\cdot\!\mathbb{E}\!\left\{\mathsf{x}^{4}_{\text{Re}}\right\}\!+\!2\!\cdot\!\left(\mathbb{E}\!\left\{\mathsf{x}^{2}_{\text{Re}}\right\}\right)^{2}$ (48)
$=2\cdot\frac{9}{20}+2\cdot\left(\frac{1}{2}\right)^{2}=\frac{7}{5},\vskip-7.11317pt$ (49)
where we used $\mathbb{E}\{\mathsf{s}^{4}\}=a^{4}/5$ and $\mathbb{E}\{\mathsf{s}^{2}\}=a^{2}/3$ for $\mathsf{s}\sim\mathcal{U}[-a,a]$ with $a=\sqrt{3/2}$.
Appendix B
@@
$=\pi\mathrm{e}^{\gamma_{0}}\mathrm{e}^{\frac{\gamma^{2}_{2}}{4\gamma_{4}}}\!\!\!\!\!\!\!\!\!\!\int\limits_{-\gamma_{2}/(2|\gamma_{4}|)}^{\infty}\!\!\!\!\!\!\left(v+\frac{\gamma_{2}}{2|\gamma_{4}|}\right)^{q}\mathrm{e}^{-|\gamma_{4}|v^{2}}\,\mathrm{d}v\vskip-2.84544pt$ (54)
@@
Using the integrals [33, entries: 3.321.2, 3.321.4, 3.321.5], we obtain closed-form expressions
@@
$\begin{split}C_{2}&=\pi\mathrm{e}^{\gamma_{0}}\left[\frac{\gamma_{2}}{4|\gamma_{4}|^{2}}+\left(\frac{1}{4|\gamma_{4}|}+\frac{\gamma_{2}^{2}}{8|\gamma_{4}|^{2}}\right)\right.\\ &\quad\qquad\quad\cdot\left.\sqrt{\frac{\pi}{|\gamma_{4}|}}\mathrm{erfc}\!\left(\frac{-\gamma_{2}}{2\sqrt{|\gamma_{4}|}}\right)\mathrm{e}^{\frac{\gamma_{2}^{2}}{4|\gamma_{4}|}}\right].\end{split}\vskip-4.26773pt$ (57)
Using $C_{0}=1$ (Tab. II), we substitute the common term in (55) into (56) and (57) yielding
@@
$C_{2}$ $=\frac{\gamma_{2}}{4|\gamma_{4}|^{2}}\left(\pi\mathrm{e}^{\gamma_{0}}+\gamma_{2}\right)+\frac{1}{2|\gamma_{4}|}.\vskip-2.84544pt$ (59)